If a gas in a closed container is pressurized from 15.0 atm to 16.0 atm and its original temperature was 25.0 °C, what would the final temperature of the gas be (in °C)? **Hint: You still need to convert to Kelvin, then convert back to Celsius also round the answer to the correct sig figs


1
Expert's answer
2021-08-17T14:37:03-0400

P1/T1 = P2/T2

T1 = 25.0 + 273 = 298.0 K

15.0/298.0 = 16.0/T2

T2 = 318.1 K